1) Villanova, Inc., has done a cost analysis for its production of rubber stamps. The following activities and cost drivers have been developed:

Activity Cost Formula
Design $5,000 + $0.05 per machine hour
Machining $25,000 + $0.01 per machine hour
Setups $35 per batch
Purchasing $50 + $15 per purchase order


Following are the actual costs of producing 35,000 rubber stamps: 1,000 machine hours; 5 batches; 30 purchase orders

Design $5,080
Machining ?
Setups ?
Purchasing $600


The following variances were given in the activity performance report:

Design ?
Machining $40 F
Setups $15 F
Purchasing ?


What is the activity variance for design?

a.$100 U

b.$40 F

c.$30 U

d.$15 F

2)

Selling and administrative costs are classified as

a.conversion costs.

b.factory overhead.

c.product costs.

d.period costs.

3)

The Davidson Company uses a weighted average process costing system. The following information was reported for the Assembly Process for January. Materials are added at the beginning of the process and are 100%.

Units: units % complete for conversions
work in process, 1/1 60,000 15%
started 105,000
work in process, 1/31 40,000 20%
Costs:
Materials Conversion
beginning work in process $ 16,500 $ 33,250
current costs $643,500 $332,500


What is the cost per unit for materials?

a.$3.90

b.$2.75

c.$4.96

d.$4.00

e.None of these choices are correct.
Answer #1

Solution 1:

Standard cost of design at 1000 machine hours = $5,000 + 1000*$0.05 = $5,050

Actual cost of design = $5,080

Activity variance for design = Standard cost - Actual cost - $5,050 - $5,080 = $30 U

Hence option c is correct.

Solution 2:

Selling and administrative costs are classified as "Period costs"

Hence option d is correct.

Solution 3:

Equivalent units of material = Units completed and transferred out + Units in ending WIP

= (60000 + 105000 - 40000) + 40000 = 165000 units

Total cost to be accounted for material = Beginning WIP + current cost = $16,500 + $643,500 = $660,000

Cost per equivalent unit of material = Total cost to account for / Equivalent unit of material

= $660,000 / 165000 = $4 per unit

Hence option d is correct,